OpenStudy (anonymous):

A population of 900 bears decreases by 15% per year. How many bears will there be in the population after 7 years? Round your answer to the nearest whole number. a. 289 b. 2,394 c. 9 d. 795 17. What is the value of 12x^-3y^-1 for x=-1 and y=5? a. -12/5 b. -180 c. 7,500 d. -5/12

OpenStudy (anonymous):

@Phebe

OpenStudy (jdoe0001):

What is the value of 12x^-3y^-1 for x=-1 and y=5? <---- that's pretty self-explanatory I'd think

OpenStudy (phebe):

im here

OpenStudy (phebe):

well @jdoe0001 is helping u so ur good :)

OpenStudy (jdoe0001):

for the 1st one, "bear" (pun intended), in mind that is a geometric sequence on that case, I'd use the compound interest formula, which is really just a geometric sequence \(\bf A=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ A=\textit{current amount}\\ P=\textit{original amount}\\ r=rate\to r\%\to \frac{r}{100}\\ n=\textit{times it compounds per year}\\ t=years\to t\) the bears are decreasing, so the rate is "negative" is happening yearly, so the cycle or period is 1

OpenStudy (jdoe0001):

hmmm hold gthe mayo \(\bf \large A=P\left(1+\frac{{\color{brown}{ -r}}}{{\color{brown}{ 1}}}\right)^{{\color{brown}{ 1}}t}\)